. The design capacity for Bakery Company is 1000 loafs of bread per day. The effective capacity is 800 loafs of bread per day and the actual output is 650 loafs of bread per day. The company goal is to produce 950 loafs of bread per day. Calculate the utilization and efficiency of the operation. If the efficiency for next month is expected to be 85%, what is the expected output?
1
Expert's answer
2020-06-09T16:44:17-0400

Since the actual level of production is 650 loaves of bread per day, the use will be 65%


"\frac{650}{1000}100=65"

The effectiveness of the operation will be 81.25%


"\frac{650}{800}100=81.25"

Next month will produce 680 loaves of bread per day


"\frac{85 \times 800}{100}=680"